OpenStudy (anonymous):

A huricane is 300 miles offshore and moving at 50 mph. For what range of time will the hurricane be between 50 and 150 miles offshore. Write the lower and upper limits separated by a comma. ex. (1,6)

OpenStudy (anonymous):

t=distance/speed t1=250/50 (50 miles offshore) t2=150/50 (150miles offshore)